# Realization of broken inversion symmetry in the charge density wave phase in EuAl4

### v1

@@ II.1 Crystal growth
High-quality single crystals of EuAl4 were grown by the self-flux method according to the procedure described by Nakamura et al. [33], using growth parameters different from our previous work [21]. Europium (Smart-elements, 99.99% purity) and aluminum (Alfa-Aesar, 99.9995%) were mixed into an alumina crucible in the elemental ratio , and then sealed under vacuum in a quartz-glass tube. The quartz-glass tube was heated to 1173 K and maintained at this temperature for two days. It was then slowly cooled down to 673 K over 110 hours. During the cooling phase the oven was slightly tilted, in order to facilitate the separation of excess Al-flux from the crystals. After cooling to room temperature, a single crystal was obtained of about mm in size (Fig. 2).
@@ III Symmetry and crystal structure
The incommensurately modulated crystal structure of EuAl4 in its CDW state is described within the superspace approach [42]. In principle, any superspace group is a candidate symmetry for the CDW phase, that is based on a modulation wave vector along and a basic-structure space group that is a ”translationsgleiche” subgroup of . A total of superspace groups exist with these properties [28]. Previous analyses of SXRD data of up to first-order satellites have shown that six of these superspace groups allow for a structure model that may describe the SXRD data well [21, 22, 28]. Present refinements confirm this finding for the SXRD-30 data (Tables S2 and S3 in [38]). These symmetries comprise two acentric tetragonal superspace groups, two acentric orthorhombic superspace groups and two centrosymmetric orthorhombic superspace groups (Fig. 4 and Table 2).
@@ IV Discussion
A second point to notice is that the two centrosymmetric orthorhombic superspace groups as well as the acentric tetragonal superspace groups are not subgroups of any superspace group based on . On the other hand, is a subgroup of both and, previously proposed, (Fig. 4). Indeed, the present acentric structure model can be considered as a distortion of the model. Major effect is the unequal displacement modulations of the Al1a and Al1b atoms (Fig. 5, and Table S7 and Fig. S2 in [38]).
@@ V Conclusions
The CDW is found to reside on the Al–Al network, in agreement with previous models. The acentric nature of the CDW modulation should be considered, when considering models for the stabilization of the skyrmion states of EuAl4.

### v2

# Broken inversion symmetry in the charge density wave phase in EuAl4

@@ II.1 Crystal growth
High-quality single crystals of EuAl4 were grown by the self-flux method according to the procedure described by Nakamura et al. [33], using growth parameters different from our previous work [21]. Europium (Smart-elements, 99.99% purity) and aluminum (Alfa-Aesar, 99.9995%) were mixed into an alumina crucible in the elemental ratio , and then sealed under vacuum in a quartz-glass tube. The quartz-glass tube was heated to 1173 K and maintained at this temperature for two days. It was then slowly cooled down to 673 K over 110 hours. The oven was tilted by 9 deg, with the materials at the lower side, thus keeping all materials together during the reaction. After cooling to room temperature, a single crystal was obtained of about mm in size (Fig. 2).
@@ III Symmetry and crystal structure
The incommensurately modulated crystal structure of EuAl4 in its CDW state is described within the superspace approach [42, 43]. In principle, any superspace group is a candidate symmetry for the CDW phase, that is based on a modulation wave vector along and a basic-structure space group that is a ”translationsgleiche” subgroup of . A total of superspace groups exist with these properties [28]. Previous analyses of SXRD data of up to first-order satellites have shown that six of these superspace groups allow for a structure model that may describe the SXRD data well [21, 22, 28]. Present refinements confirm this finding for the SXRD-30 data (Tables S2 and S3 in [38]). These symmetries comprise two acentric tetragonal superspace groups, two acentric orthorhombic superspace groups and two centrosymmetric orthorhombic superspace groups (Fig. 4 and Table 2).
@@ IV Discussion
A second point to notice is that the two centrosymmetric orthorhombic superspace groups as well as the acentric tetragonal superspace groups are not subgroups of any superspace group based on . On the other hand, is a subgroup of both and, previously proposed, (Fig. 4). Indeed, the present acentric structure model can be considered as a distortion of the model (Compare the discussion in Section S6 in the Supplemental Material [38]). Initially, it appears that all atoms exhibit similar displacements along (Figs. 5 and 6(a), and Table S7 and Figs. S1–S3 in [38]). However, major effect is the unequal displacement modulations of the Al1a and Al1b atoms into the direction of (Figs. 5 and 6(b), and Table S7 and Figs. S1–S3 in [38]).
@@ V Conclusions
The CDW is found to reside on the Al–Al network, in agreement with previous models. The acentric nature of the CDW modulation is most clearly established by the unequal modulations of the Al1a and Al1b atoms. It should be considered, when considering models for the stabilization of the skyrmion states of EuAl4.
@@ Acknowledgements.
- Janssen et al. [2018] T. Janssen, G. Chapuis, and M. de Boissieu, Aperiodic Crystals. From Modulated Phases to Quasicrystals, 2nd ed. (Oxford University Press, Oxford, 2018).